# NMPCB: A Lightweight and Safety-Critical Motion Control Framework

### v1

@@ III-D Framework Overview
The framework consists of a planner and a controller. The planner utilizes Neural Dubins Model to determine the next target point and generates a Dubins curve from the current point to the target point, which serves as a reference trajectory and is passed to the control module. The control module solves equation (23) to output the optimal control commands, ensuring that the robot moves towards the goal without collisions.
We focus on the motion control problem of mobile robots, hence adopting the bicycle model as the kinematic model.The state vector of the robot $\mathbf{x}$ is defined as $[x,y,v,\theta]$, where $x$ and $y$ represent the coordinates of the rear axle center, $v$ is the velocity, and $\theta$ is the yaw angle. The control vector $\mathbf{u}$ is defined as $[a,\delta]$, where $a$ is the acceleration and $\delta$ is the steering angle of the front wheel. The kinematic equations are given by:
@@
$\theta_{k+1}=\theta_{k}+\frac{v_{k}\tan{\delta_{k}}}{L}\Delta t,$ (25d)
where $\Delta t$ represents the time step and $L$ denotes the wheelbase.
@@
$q(\mathbf{x}_{t+k|t},\mathbf{u}_{t+k|t})$ $=\|\mathbf{x}_{t+k|t}-\bar{\mathbf{x}}_{t+k|t}\|^{2}_{Q}+\|\mathbf{u}_{t+k|t}\|^{2}$
@@
For the point mass model of the robotic, we can ensure effective obstacle avoidance by setting a safety distance $D_{safe}$, which also alleviates the issue of decreased solution speed due to the incorporation of distant obstacles in the optimization problem.
@@
In the following section, we will conduct comparative experiments of the aforementioned algorithms across various scenarios.
@@ IV Experiments and Results
In this section, we conduct numerical simulations of the NMPCB framework and perform ablation experiments to verify that both the planner and the controller significantly enhance the motion control problem. Furthermore, we have conducted real-world experiments with the proposed framework to validate its performance in actual environments.
(a) Dubins Curve + MPC-DCBF
(b) Dubins Curve + MDD-I
@@
Figure 4: Comparison of trajectories among different algorithms in numerical simulation experiments.
IV-A Datasets Construction
Due to the scarcity of high-quality path planning datasets for deep learning in unstructured environments, we have constructed a dataset that includes information on obstacles and the robot’s motion trajectories. The generation of motion trajectories within the dataset is facilitated by the RDA [11] algorithm. We generated a substantial amount of motion trajectory data in unstructured scenarios by randomly creating obstacles and assigning start and end points within a certain range.
@@ IV-B Numerical Simulation
We have prepared the following four combinations of motion control algorithms for comparative experiments: (1) A Dubins curve connecting the starting point to the goal point is utilized as the planner, with the MPC-DCBF serving as the controller. (2) A Dubins curve connecting the starting point to the goal point is utilized as the planner, with the MDD-I serving as the controller. (3) The Neural Dubins Model is utilized as the planner, with the MDD-I serving as the controller. (4) The Neural Dubins Model is utilized as the planner, with the MDD-III serving as the controller. We have set the parameters for the aforementioned algorithms to $N=11$, $N_{{CBF}}=10$, and $\gamma=0.9$. The optimization problems are implemented in Python using CasADi[25] as the modeling language and solved with IPOPT [26] on Ubuntu 18.04.
@@
We validated our proposed framework in intelligent robot simulator (ir-sim) [^1] https://github.com/hanruihua/ir_sim, which is a Python-based simulator for robotic algorithm development. We represent obstacles with black cubes, denote the area traversed by the robot with a yellow border, and illustrate the reference trajectory provided by the planner to the controller with colored lines.
In Fig. 4LABEL:sub@fig4a, the red dots indicate points where the solver failed to find a solution, demonstrating the poor performance of MPC-DCBF in complex scenarios. Fig. 4LABEL:sub@fig4b shows that MDD-I can adapt to complex environments, however, due to the Dubins curve not accounting for obstacles, the controller collided while following the reference trajectory. Fig. 4LABEL:sub@fig4c - 4LABEL:sub@fig4d illustrate that, upon employing the Neural Dubins Model, all two algorithms are capable of planning a suitable collision-free trajectory. This demonstrates that the Neural Dubins Model provides an appropriate reference path. However, there is a significant difference in computation time among the two algorithms, especially in densely obstacle-populated areas, where the computation time for MDD-II often reaches around 0.15 seconds, whereas the computation time for MDD-I does not exceed 0.1 seconds. Fig. 5 reflects the disparity in their computation times.
Figure 5: Comparison of computation cost of MDD-I and MDD-II. The horizontal axis represents the time step, and the vertical axis represents the solution time, with the unit being seconds.
To better compare the differences in success rate and average solution time among the two algorithms, we conducted comparative experiments across 50 randomly generated scenarios, and the results are presented in Table 1.
TABLE I: Performance of MDD across 50 random scenarios
Success Rate Average Solving Time Average Max Solving Time
Neural Dubins MDD-I 0.86 Four 0.172
Neural Dubins MDD-II 0.82 Four 0.851
Here, the success rate refers to the ratio of the number of scenarios where the robot successfully reaches the destination without collision to the total number of scenarios. The average solving time is the average single-step solving time in successful scenarios. The average maximum solving time refers to the average of the maximum single-step solving time in each successful scenario.
(a) Dubins Curve + MPC-DCBF
(b) Dubins Curve + MDD-I
(c) Neural Dubins Model + MDD-I
(d) Neural Dubins Model + MDD-II
@@
Figure 6: Comparison of trajectories among different algorithms in real-world experiments. Subfigures (c) and (d) correspond to control module frequencies of 10 Hz. The red arrow indicates the location where a collision occurred.
It can be observed that MDD-I and MDD-II exhibit similar success rates. However, both average solving time and average maximum solving time for MDD-I are significantly lower than those for MDD-II.
Figure 7: The Ackermann robot.
@@ IV-C Real-World Experiments
Utilizing the aforementioned hardware, we conducted comparative experiments of the four algorithms from numerical simulations in real-world settings. In Fig. 6LABEL:sub@fig6a, the MPC-DCBF algorithm faced persistent solution failures in the presence of multiple obstacles, ultimately resulting in a collision. In Fig. 6LABEL:sub@fig6b, although no solution failures occurred, collisions still took place due to the use of the Dubins curve as the reference trajectory. Comparing Fig. 6LABEL:sub@fig6c and Fig. 6LABEL:sub@fig6d, the MDD-II algorithm resulted in a collision due to the robot’s control module frequency exceeding the solution frequency of MDD-II. However, the MDD-I algorithm demonstrated superior real-time performance, successfully reaching the destination.
V CONCLUSIONS

### v2

@@ III-D Framework Overview
The framework consists of a planner and a controller. The planner utilizes Neural Dubins Model to determine the next target point and generates a Dubins curve from the current point to the predicted target point, which serves as a reference trajectory and is passed to the control module. The control module solves optimization problem (23) to output the optimal control commands, ensuring that the robot moves towards the goal without collisions.
We focus on the motion control problem of mobile robots, hence adopting the bicycle model as the kinematic model. The state vector of the robot $\mathbf{x}$ is defined as $x,y,v,\theta]$, where $x$ and $y$ represent the coordinates of the rear axle center, $v$ is the velocity, and $\theta$ is the yaw angle. The control vector $\mathbf{u}$ is defined as $a,\delta]$, where $a$ is the acceleration and $\delta$ is the steering angle of the front wheel. The kinematic equations are given by:
@@
$\theta_{k+1}=\theta_{k}+\frac{v_{k}\tan{(\delta_{k})}}{L}\Delta t,$ (25d)
where $\Delta t$ denotes the time step and $L$ denotes the wheelbase.
@@
$q(\mathbf{x}_{t+k|t},\mathbf{u}_{t+k|t})$ $=\|\mathbf{x}_{t+k|t}-\bar{\mathbf{x}}_{t+k|t}\|^{2}+\|\mathbf{u}_{t+k|t}\|^{2}$
@@
For the point mass model of the robot, we can ensure effective obstacle avoidance by setting a safety distance $D_{\text{safe}}$, which also alleviates the issue of decreased solution speed due to the incorporation of distant obstacles in the optimization problem.
@@ IV Experiments and Results
In this section, we perform numerical simulations of the NMPCB framework and conduct repeated experiments across diverse scenarios to verify that the framework can significantly improve the solution speed of the motion control problem while ensuring robot safety. Furthermore, we have conducted real-world experiments with the proposed framework to validate its performance in actual environments.
(a) RDA Planner + MPC-DCBF
(b) RDA Planner + MDD-I
@@
(e) RDA Planner + MPC-DCBF
(f) RDA Planner + MDD-I
(g) Neural Dubins Model + MDD-I
(h) Neural Dubins Model + MDD-II
Figure 4: Comparison of trajectories among different algorithms in numerical simulations.
IV-A Datasets Construction and Training Details
Due to the scarcity of high-quality path planning datasets for deep learning in unstructured environments, we have constructed a dataset that includes information on obstacles and the robot’s motion trajectories. The generation of motion trajectories within the dataset is facilitated by the RDA Planner [11]. The simulation environment is configured as a 50 m $\times$ 50 m square workspace. Start and goal poses are stochastically sampled within bounded regions located at the square’s opposite corners. The obstacle set comprises four to six obstacles. A subset is uniformly sampled across the entirety of the 50 m $\times$ 50 m workspace, while the remaining obstacles are stochastically positioned within a narrow corridor centered on the arena’s diagonal axis. Obstacle dimensions are likewise randomized within a bounded interval, ensuring that neither excessive size nor diminutive scale compromises dataset fidelity or task validity. Subsequently, the planner is employed to generate collision-free trajectories. The obstacle configurations and corresponding trajectory data are then aggregated to constitute a single training instance.
@@ IV-A Datasets Construction and Training Details
Training was performed on a workstation equipped with an NVIDIA RTX 4060 GPU and 16 GB of system memory. The learning rate was set to 0.001, and the Adam optimizer was adopted to accelerate convergence. A batch size of 32 was employed to balance memory utilization and computational efficiency. Finally, the model was trained for 200 epochs to ensure full convergence.
@@ IV-B Numerical Simulation
We have prepared the following four combinations of motion control algorithms for comparative experiments: (1) The RDA is utilized as the planner, with the MPC-DCBF serving as the controller. (2) The RDA is utilized as the planner, with the MDD-I serving as the controller. (3) The Neural Dubins Model is utilized as the planner, with the MDD-I serving as the controller. (4) The Neural Dubins Model is utilized as the planner, with the MDD-II serving as the controller. We have set the parameters for the aforementioned algorithms to $N=11$, $N_{{\text{CBF}}}=10$, and $\gamma=0.9$. The optimization problems are implemented in Python using CasADi[26] as the modeling language and solved with IPOPT [27] on Ubuntu 18.04.
When the degree of constraint violation of a numerical solution is less than the feasibility tolerance of the solver, the solver will regard this solution as a feasible point. Consequently, violations of the obstacle avoidance constraint may occur, leading to collisions.
@@
We validated our proposed framework in intelligent robot simulator (ir-sim) [^1] https://github.com/hanruihua/ir_sim, which is a Python-based simulator for robotic algorithm development. We represent obstacles with black polygons, denote the area traversed by the robot with a yellow border, and illustrate the reference trajectory provided by the planner to the controller with colored lines. Comparative simulations were conducted in both square scenarios and line scenarios to comprehensively evaluate the proposed framework.
(a) Square Scenario
(b) Line Scenario
Figure 5: Comparison of computation cost of three algorithms. It illustrates the deviation of the computational cost of three algorithms at each time step from the $0.1$s benchmark line in Fig. 4.
(a) RDA Planner + MDD-I
(b) Neural Dubins Model + MDD-I
(c) Neural Dubins Model + MDD-II
(d) Line Scenario with 4 Obstacles
(e) Square Scenario with 6 Obstacles
(f) Square Scenario with 7 Obstacles
(g) Square Scenario with 8 Obstacles
(h) Line Scenario with 6 Obstacles
@@
Figure 6: Comparison of trajectories among different algorithms in real-world experiments. During the experiments, the control module frequency was set to 10 Hz.
In Fig. 4LABEL:sub@fig4a and Fig. 4LABEL:sub@fig4e, the red dots denote locations where the solver fails to obtain a feasible solution, demonstrating the limited adaptability of MPC-DCBF in complex scenarios. Compared with Fig. 4LABEL:sub@fig4b and Fig. 4LABEL:sub@fig4f, the MDD-I controller demonstrates superior adaptability to complex environments, exhibiting no solver failures. Similarly, inspection of Fig. 4LABEL:sub@fig4c - LABEL:sub@fig4d and Fig. 4LABEL:sub@fig4g - LABEL:sub@fig4h reveals that, upon integrating the Neural Dubins Model, both frameworks consistently generate collision-free trajectories that satisfy all imposed constraints. Nevertheless, the three frameworks exhibit markedly disparate computational overheads. In obstacle-dense regions, the average solution times for RDA Planner with MDD-I and Neural Dubins Model with MDD-II cluster around 0.15 seconds, whereas Neural Dubins Model with MDD-I consistently solves within 0.10 seconds. These quantitative discrepancies are further illustrated in Fig. 5.
To quantitatively and visually assess the disparities in success rates and average solution times among the algorithms, we conducted comparative experiments across 50 randomly generated square scenarios and 50 randomly generated line scenarios. The aggregated results are presented in Table I and Table II.
Here, the success rate refers to the ratio of the number of scenarios where the robot successfully reaches the destination without collision to the total number of scenarios. Across all successful scenarios, we separately report the single-step average runtime for both the planner and the controller, and additionally computed the mean of the maximum single-step execution times.
TABLE I: Performance across 50 random square scenarios
Success Rate
Planner
Controller
Max
RDA
MPC-DCBF
0.16 - - -
RDA
MDD-I
0.88 54ms 32ms 167ms
Neural Dubins
MDD-I
0.86 5ms 33ms 86ms
Neural Dubins
MDD-II
0.88 5ms 84ms 181ms
TABLE II: Performance across 50 random line scenarios
Success Rate
Planner
Controller
Max
RDA
MPC-DCBF
0.12 - - -
RDA
MDD-I
0.76 63ms 44ms 176ms
Neural Dubins
MDD-I
0.76 5ms 48ms 92ms
Neural Dubins
MDD-II
0.80 5ms 93ms 193ms
Owing to the prohibitively low success rate of RDA Planner with MPC-DCBF, its runtime statistics lack statistical reliability and are therefore excluded from the reported results. The remaining three algorithms exhibit comparable success rates. However, the Neural Dubins Model markedly outperforms the RDA Planner in computational efficiency. Moreover, MDD-I yields a significantly lower single-step solution time than MDD-II. Collectively, the Neural Dubins Model paired with MDD-I demonstrates a pronounced advantage in overall solution speed.
@@ IV-C Real-World Experiments
Figure 7: The Ackermann mobile robot.
Utilizing the aforementioned hardware, we carried out real-world experimental validation. Inspection of Fig. 6LABEL:sub@fig6a - LABEL:sub@fig6c reveals that, because the control module’s execution frequency exceeds the algorithm’s solution frequency, neither RDA Planner with MDD-I nor Neural Dubins Model with MDD-II is able to reach the goal. In contrast, Neural Dubins Model with MDD-I demonstrates superior real-time performance and successfully reaches the goal. Fig. 6LABEL:sub@fig6d - LABEL:sub@fig6h show that the Neural Dubins Model with MDD-I was further evaluated under a diverse set of challenging scenarios. In every trial the robot successfully reached the goal, thereby confirming the algorithm’s robust feasibility in complex environments.
V Conclusion